# TextOCVP: Object-Centric Video Prediction with Language Guidance

### v2

@@ Abstract
Understanding and forecasting future scene states is critical for autonomous agents to plan and act effectively in complex environments. Object-centric models, with structured latent spaces, have shown promise in modeling object dynamics and predicting future scene states, but often struggle to scale beyond simple synthetic datasets and to integrate external guidance, limiting their applicability in robotics. To address these limitations, we propose TextOCVP, an object-centric model for video prediction guided by textual descriptions. TextOCVP parses an observed scene into object representations, called slots, and utilizes a text-conditioned transformer predictor to forecast future object states and video frames. Our approach jointly models object dynamics and interactions while incorporating textual guidance, enabling accurate and controllable predictions. TextOCVP’s structured latent space offers a more precise control of the forecasting process, outperforming several video prediction baselines on two datasets. Additionally, we show that structured object-centric representations provide superior robustness to novel scene configurations, as well as improved controllability and interpretability, enabling more precise and understandable predictions. Videos and code are available at https://play-slot.github.io/TextOCVP.
@@ 1 Introduction
Understanding and reasoning about the environment is essential for enabling autonomous systems to better comprehend their surroundings, predict future events, and adapt their actions accordingly. Humans achieve these capabilities by perceiving the environment as a structured composition of individual objects that interact and evolve dynamically over time (Kahneman, Treisman, and Gibbs 1992). Neural networks equipped with such compositional inductive biases have shown the ability to learn structured object-centric representations of the world, which enable desirable properties, such as out-of-distribution generalization (Dittadi et al. 2022), compositionality (Greff, Van Steenkiste, and Schmidhuber 2020), or sample efficiency (Mosbach et al. 2025).
Recent advances in unsupervised object-centric representation learning have progressed from extracting object representations in synthetic images (Locatello et al. 2020; Lin et al. 2020) to modeling objects in video (Kipf et al. 2022; Singh, Wu, and Ahn 2022) and scaling to real-world scenes (Seitzer et al. 2023; Zadaianchuk, Seitzer, and Martius 2024). These developments have enabled object-level dynamics modeling for future prediction and planning. Notably, approaches like SlotFormer (Wu et al. 2023) or OCVP (Villar-Corrales, Wahdan, and Behnke 2023) introduced object-centric prediction models that explicitly model spatio-temporal relationships between objects, shifting away from image-level approaches that ignore scene compositionality. Despite these advancements, current object-centric methods struggle with complex object appearances and dynamics, and lack mechanisms to incorporate external guidance, thus limiting their scalability and broader applicability.
Refer to captiona) Object-Centric ParsingLearned ObjectsRefer to captionb) Text-Conditioned Video Prediction
‘Put the green block
in the cyan bowl.’
‘Put the green block
in the red bowl.’
@@
We evaluate our approach for text-conditioned video prediction through extensive experiments on two distinct datasets. Our results show that TextOCVP outperforms existing text-conditioned video prediction methods by effectively leveraging structured object-centric representations, particularly on scenes featuring multiple moving objects.
To verify the effectiveness of object-centric representations for text-guided video prediction, we conduct an in-depth analysis of TextOCVP. Specifically, we demonstrate that its structured latent space enables accurate and controllable video prediction by aligning language instructions with the corresponding objects, outperforming baselines that rely on holistic scene representations. Beyond generation quality, TextOCVP offers improved interpretability and exhibits strong robustness to novel scene configurations, including varying numbers of objects or previously unseen colors.
@@
  • Through extensive evaluations, we show that TextOCVP outperforms existing text-conditioned video prediction models by leveraging object-centric representations.
@@ 2.1 Object-Centric Learning
Representation learning, the ability to extract meaningful features from data, often improves model performance by enhancing its understanding of the input space (Bengio, Courville, and Vincent 2013). Object-centric representation methods aim to parse an image or video into a set of object components in an unsupervised manner. These objects are typically represented as unconstrained embeddings (called slots) (Locatello et al. 2020; Kipf et al. 2022; Singh, Wu, and Ahn 2022), patch-based representations (Lin et al. 2020), factored latent vectors (Greff et al. 2019), or explicit object prototypes (Villar-Corrales and Behnke 2022; Monnier, Groueix, and Aubry 2020). These methods have demonstrated promise in learning object representations across diverse domains, ranging from synthetic images (Locatello et al. 2020; Lin et al. 2020) to videos (Kipf et al. 2022; Singh, Wu, and Ahn 2022; Elsayed et al. 2022), and real-world scenes (Seitzer et al. 2023; Zadaianchuk, Seitzer, and Martius 2024). The learned object representations benefit downstream tasks, such as reinforcement learning (Mosbach et al. 2025) or visual-question answering (Wu et al. 2023).
@@ 2.2 Video Prediction
Video prediction (VP) is the task of forecasting the upcoming $T$ video frames conditioned on the preceding $C$ seed frames (Oprea et al. 2020). Several methods have been proposed to address this challenge, leveraging 2D convolutions (Gao et al. 2022; Chiu, Adeli, and Niebles 2020), 3D convolutions (Tulyakov et al. 2018), recurrent neural networks (RNNs) (Denton and Fergus 2018; Villar-Corrales et al. 2022; Wang et al. 2022), Transformers (Rakhimov et al. 2021; Ye and Bilodeau 2022), or diffusion models (Höppe et al. 2022; Ho et al. 2022).
Object-Centric Video Prediction
@@
Object-centric VP presents a structured approach that explicitly models the dynamics and interactions of individual objects to forecast future video frames. These methods typically involve three main steps: parsing seed frames into object representations, forecasting future object states using a dynamics model, and rendering video frames from the predicted object representations. Various approaches have addressed this task using different architectural priors, such as RNNs (Creswell et al. 2021; Nguyen et al. 2024) or transformers (Wu, Yoon, and Ahn 2021; Wu et al. 2023; Villar-Corrales, Wahdan, and Behnke 2023). Despite promising results, these models are limited to simple deterministic datasets or rely on action-conditioning (Mosbach et al. 2025; Villar-Corrales and Behnke 2025). In contrast, our model forecasts future frames conditioned on past object slots and text descriptions.
Text-Conditioned Video Prediction
@@
Text-conditioned VP models leverage text descriptions to provide appearance, motion and action cues that guide the generation of future frames. This task was first proposed by Hu, Luo, and Chen (2022), who utilized a VQ-VAE to encode images into visual token representations, and modeled the scene dynamics with an axial transformer to jointly process visual tokens with text descriptions. Similarly, approaches like TVP (Song et al. 2024) and MMVG (Fu et al. 2023) address this task using RNNs or masked transformers, respectively. More recently, several methods leverage diffusion models for text-guided VP (Blattmann et al. 2023; Ni et al. 2023; Chen et al. 2024; Xing et al. 2024; Chen et al. 2023; Gu et al. 2024). These approaches encode video frames into discrete token sequences via pretrained quantized autoencoders, and leverage pretrained language models to guide the diffusion-based generation with text features. To further improve temporal coherence and semantic alignment, several diffusion-based methods leverage specialized conditioning strategies (Xing et al. 2024; Chen et al. 2023) or attention mechanisms (Gu et al. 2024), among others. While effective, these models operate on holistic or spatial representations, and require large-scale compute and data. In contrast, TextOCVP adopts a more structured and efficient approach, explicitly modeling object dynamics using slot representations, where each object in the scene is represented by a distinct embedding.
@@
Concurrently with our work, Wang et al. (2025) and Jeong et al. (2025) combine object-centric learning with autoregressive diffusion models and transformers for text-guided video prediction on simple synthetic datasets. In contrast, we model object dynamics with an autoregressive transformer, evaluate on more complex robotics simulations, and perform an in-depth analysis of the properties of object-centric representations for text-conditioned video prediction.
@@ 3 Method
We propose two different TextOCVP variants, which differ in the underlying object-centric decomposition modules. Specifically, $\text{TextOCVP}_{\text{SAVi}}$ leverages SAVi (Kipf et al. 2022), whereas $\text{TextOCVP}_{\text{DINO}}$ extends the recent DINOSAUR (Seitzer et al. 2023) framework for recursive object-centric video decomposition and video rendering.
@@ 3.1 Scene Parsing
The scene parsing module decomposes a video sequence $\mathbf{X}_{1:{\tau}}$ into a set of permutation-invariant object representations called slots $\mathbf{S}_{1:\tau}=(\textbf{S}_{1},...,\textbf{S}_{\tau})$, with $\textbf{S}_{t}=(\textbf{s}_{t}^{1},...,\textbf{s}_{t}^{N_{\mathbf{S}}})$, where each slot $\textbf{s}\in{\mathbb{R}}^{D}$ represents a single object. For scene parsing, we adopt the recursive object-centric video decomposition framework from Kipf et al. (2022).
At time step $t$, the corresponding input frame $\textbf{X}_{t}$ is encoded with a feature extractor module into a set of $D_{h}$-dimensional feature maps $\textbf{h}_{t}\in{\mathbb{R}}^{L\times D_{h}}$ representing $L$ spatial locations. The feature extractor is a convolutional neural network in our $\text{TextOCVP}_{\text{SAVi}}$ variant and a DINO-pretrained vision transformer (Caron et al. 2021) in $\text{TextOCVP}_{\text{DINO}}$. These feature maps are processed with Slot Attention (Locatello et al. 2020), which updates the previous slots $\textbf{S}_{t-1}$ based on visual features from the current frame following an iterative attention mechanism. Namely, Slot Attention performs cross-attention, with the attention weights normalized over the slot dimension, thus encouraging the slots to compete to represent parts of the input. It then updates the slots using a Gated Recurrent Unit (Cho et al. 2014) (GRU). Formally, Slot Attention updates the previous slots $\textbf{S}_{t-1}$ by:
@@
Figure 2: Overview of TextOCVP. TextOCVP parses the reference frame $\textbf{X}_{1}$ into object representations $\textbf{S}_{1}$. The text-conditioned object-centric predictor models object dynamics and interactions, incorporating information from the description $\mathbf{C}$ to predict future object states $\hat{\mathbf{S}}_{2:T+1}$, which can be used to render future video frames $\hat{\mathbf{X}}_{2:{T+1}}$.
@@ 3.2 Text-Conditioned Object-Centric Predictor
Our proposed text-conditioned predictor module, depicted in Fig. 3, autoregressively forecasts future object states conditioned on the object slots from the reference frame $\textbf{S}_{1}$ and a textual description $\mathcal{C}$.
To condition the prediction process, the text description $\mathcal{C}$ is encoded into text token embeddings $\mathbf{C}$ employing an encoder-only transformer module. We experiment with different variants of this module, including a vanilla transformer encoder (Vaswani et al. 2017) or a pretrained T5 (Raffel et al. 2020) module.
@@
Each layer of our predictor module mirrors the transformer decoder architecture (Vaswani et al. 2017). First, a self-attention layer enables every slot to attend to all other object representations in the sequence, modeling the spatio-temporal relations between objects. Subsequently, a text-to-slot cross-attention layer enhances the slot representations by incorporating relevant features from the text embeddings, guiding the prediction process to align with the motion and dynamics described in the textual caption. Finally, an MLP is applied for each token. This process is repeated in every predictor layer, resulting in the predicted object slots of the subsequent time step $\hat{\textbf{S}}_{t+1}$. Furthermore, we apply a residual connection from $\textbf{S}_{t}$ to $\hat{\textbf{S}}_{t+1}$, which improves the prediction temporal consistency. This process is repeated autoregressively to obtain slot predictions for $T$ subsequent time steps.
@@ 3.3 Video Rendering
The video rendering module decodes the predicted slots $\hat{\textbf{S}}_{t}$ to render the corresponding video frame $\hat{\textbf{X}}_{t}$. We leverage two variants of the video rendering module, for our $\text{TextOCVP}_{\text{SAVi}}$ and $\text{TextOCVP}_{\text{DINO}}$ variants, respectively.
@@
This variant independently decodes each slot in $\hat{\textbf{S}}_{t}$ with a CNN-based Spatial Broadcast Decoder (Watters et al. 2019), rendering an object image $\textbf{o}_{t}^{n}$ and mask $\textbf{m}_{t}^{n}$ for each slot $\textbf{s}_{t}^{n}$. The object masks are normalized across the slot dimension, and the representations are combined via a weighted sum to render video frames:
@@
Figure 3: Overview of our text-conditioned object-centric video predictor. Our predictor employs a transformer-based architecture to model the spatio-temporal dynamics of object slots via self-attention, while incorporating textual guidance through a text-to-slot cross-attention mechanism.
@@
This decoder variant decodes the object slots in two distinct stages. First, following DINOSAUR (Seitzer et al. 2023), an MLP-based Spatial Broadcast Decoder (Watters et al. 2019) is used to generate object features along with their corresponding masks. Similar to the $\text{TextOCVP}_{\text{SAVi}}$ decoder, the object masks are normalized and combined with the object features in order to reconstruct the encoded features $\hat{\textbf{h}}_{t}\in{\mathbb{R}}^{L\times D_{h}}$. In the second stage, the reconstructed features $\hat{\textbf{h}}_{t}$ are arranged into a grid format and processed with a CNN decoder to generate the corresponding video frame $\hat{\textbf{X}}_{t}$.

### v3

@@ Abstract
Understanding and forecasting future scene states is critical for autonomous agents to plan and act effectively in complex environments. Object-centric models, with structured latent spaces, have shown promise in modeling object dynamics and interactions in order to predict future scene states, but often struggle to scale beyond simple synthetic datasets and to integrate external guidance, limiting their applicability in robotic environments. To address these limitations, we propose TextOCVP, an object-centric model for video prediction guided by textual descriptions. TextOCVP parses an observed scene into object representations, called slots, and utilizes a text-conditioned transformer predictor to forecast future object states and video frames. Our approach jointly models object dynamics and interactions while incorporating textual guidance, enabling accurate and controllable predictions. TextOCVP’s structured latent space offers a more precise control of the forecasting process, outperforming several video prediction baselines on two datasets. Additionally, we show that structured object-centric representations provide superior robustness to novel scene configurations, as well as improved controllability and interpretability, enabling more precise and understandable predictions. Videos and code are available on our project website.
footnotetext: $\dagger$ denotes equal contribution.
@@ 1 Introduction
Understanding and reasoning about the environment is essential for enabling autonomous systems to better comprehend their surroundings, predict future events, and adapt their actions accordingly. Humans achieve these capabilities by perceiving the environment as a structured composition of individual objects that interact and evolve dynamically over time (Kahneman et al., 1992). Neural networks equipped with such compositional inductive biases have shown the ability to learn structured object-centric representations of the world, which enable desirable properties, such as out-of-distribution generalization (Dittadi et al., 2022), compositionality (Greff et al., 2020), or sample efficiency (Mosbach et al., 2025).
Recent advances in unsupervised object-centric representation learning have progressed from extracting object representations in synthetic images (Locatello et al., 2020; Lin et al., 2020) to modeling objects in video (Kipf et al., 2022; Singh et al., 2022) and scaling to real-world scenes (Seitzer et al., 2023). These developments have enabled object-level dynamics modeling for future prediction and planning. Notably, approaches like SlotFormer (Wu et al., 2023) or OCVP (Villar-Corrales et al., 2023) introduced object-centric prediction models that explicitly model spatio-temporal relationships between objects, shifting away from image-level approaches that ignore scene compositionality. Despite these advancements, current object-centric methods struggle with complex object appearances and dynamics, and lack mechanisms to incorporate external guidance, thus limiting their scalability and broader applicability.
[…]
@@
We evaluate our approach for text-conditioned video prediction through extensive experiments on two distinct datasets. Our results show that TextOCVP outperforms several other text-conditioned video prediction methods by effectively leveraging structured object-centric representations, particularly on scenes featuring multiple moving objects. Furthermore, we conduct an in-depth analysis of TextOCVPto verify the effectiveness of object-centric representations for text-guided video prediction. Specifically, we demonstrate that its structured latent space enables accurate and controllable video prediction by aligning language instructions with the corresponding objects, outperforming baselines that rely on holistic scene representations. Beyond generation quality, TextOCVP offers improved interpretability and exhibits strong robustness to novel scene configurations, including varying numbers of objects or previously unseen colors.
@@
  • Through extensive evaluations, we show that TextOCVP outperforms other existing text-conditioned video prediction models by leveraging object-centric representations.
@@ 2.1 Object-Centric Learning
Representation learning—the ability to extract meaningful features from data—often improves model performance by enhancing its understanding of the input space (Bengio et al., 2013). Object-centric representation methods aim to parse an image or video into a set of object components in an unsupervised manner. These objects are typically represented as unconstrained embeddings (called slots) (Locatello et al., 2020; Kipf et al., 2022; Singh et al., 2022), patch-based representations (Lin et al., 2020), factored latent vectors (Greff et al., 2019), or object prototypes (Villar-Corrales and Behnke, 2022; Monnier et al., 2020). These methods have demonstrated promise in learning object representations across diverse domains, ranging from synthetic images (Locatello et al., 2020; Lin et al., 2020) to videos (Kipf et al., 2022; Singh et al., 2022), and real-world scenes (Seitzer et al., 2023; Zadaianchuk et al., 2024). The learned object representations benefit downstream tasks, such as reinforcement learning (Mosbach et al., 2025) or visual-question answering (Wu et al., 2023).
@@ 2.2 Video Prediction
Video prediction (VP) is the task of forecasting the upcoming $T$ video frames conditioned on the preceding $C$ seed frames (Oprea et al., 2020). Several methods have been proposed to address this task, using 2D convolutions (Gao et al., 2022; Chiu et al., 2020), 3D convolutions (Tulyakov et al., 2018), recurrent neural networks (RNNs) (Denton and Fergus, 2018; Villar-Corrales et al., 2022; Wang et al., 2022), transformers (Rakhimov et al., 2021; Ye and Bilodeau, 2022), or diffusion models (Höppe et al., 2022; Ho et al., 2022).
2.2.1 Object-Centric Video Prediction
@@
Object-centric VP presents a structured approach that explicitly models the dynamics and interactions of individual objects to forecast future video frames. These methods typically involve three main steps: parsing seed frames into object representations, predicting future object states using a dynamics model, and rendering video frames from the predicted object representations. Various approaches have addressed this task using different architectural priors, such as RNNs (Creswell et al., 2021; Nguyen et al., 2024) or transformers (Wu et al., 2021; 2023; Villar-Corrales et al., 2023; Daniel and Tamar, 2024). Despite promising results, these models are limited to simple deterministic datasets or rely on action-conditioning (Mosbach et al., 2025; Villar-Corrales and Behnke, 2025). In contrast, our model forecasts future frames conditioned on past object slots and text descriptions.
2.2.2 Text-Conditioned Video Prediction
@@
Text-conditioned VP models leverage text descriptions to provide appearance, motion and action cues that guide the generation of future frames. This task was first proposed by Hu et al. (2022), who utilized a VQ-VAE to encode images into visual token representations, and modeled the scene dynamics with an axial transformer to jointly process visual tokens with text descriptions. Similarly, approaches like TVP (Song et al., 2024) and MMVG (Fu et al., 2023) address this task using RNNs or masked transformers, respectively. More recently, several methods leverage diffusion models for text-guided VP (Blattmann et al., 2023; Ni et al., 2023; Chen et al., 2024; Xing et al., 2024; Chen et al., 2023; Gu et al., 2024). These approaches encode video frames into discrete token sequences via pretrained quantized autoencoders, and leverage pretrained language models to guide the diffusion-based generation with text features. To further improve temporal coherence and semantic alignment, several diffusion-based methods leverage specialized conditioning strategies (Xing et al., 2024; Chen et al., 2023) or attention mechanisms (Gu et al., 2024), among others. While effective, these models operate on holistic or spatial representations, and require large-scale compute and data. In contrast, TextOCVP adopts a more structured and efficient approach, explicitly modeling object dynamics using slot representations, where each object in the scene is represented by a distinct embedding.
@@
Concurrently with our work, Wang et al. (2025) and Jeong et al. (2025) also combine object-centric learning with autoregressive diffusion models and transformers for text-guided video prediction on simple synthetic datasets. In contrast, we evaluate on more complex robotic simulations, and perform an in-depth analysis of the properties of object-centric representations for text-conditioned video prediction.
@@ 3 Method
We propose two different TextOCVP variants, which differ in the underlying object-centric decomposition modules. Specifically, $\text{TextOCVP}_{\text{SAVi}}$ leverages SAVi (Kipf et al., 2022), whereas $\text{TextOCVP}_{\text{DINO}}$ extends DINOSAUR (Seitzer et al., 2023) for recursive object-centric video decomposition and video rendering.
[…]
Figure 2: Overview of TextOCVP. Our model parses the reference frame $\textbf{X}_{1}$ into object representations $\textbf{S}_{1}$. The text-conditioned object-centric predictor models object dynamics, incorporating information from the description $\mathbf{C}$ to predict future object states $\hat{\mathbf{S}}_{2:T+1}$, which can be decoded into video frames $\hat{\mathbf{X}}_{2:{T+1}}$.
@@ 3.1 Scene Parsing
The scene parsing module decomposes a video sequence $\mathbf{X}_{1:{\tau}}$ into a set of permutation-invariant object representations called slots $\mathbf{S}_{1:\tau}=(\textbf{S}_{1},\ldots,\textbf{S}_{\tau})$, with $\textbf{S}_{t}=(\textbf{s}_{t}^{1},\ldots,\textbf{s}_{t}^{N_{\mathbf{S}}})$, where each slot $\textbf{s}\in{\mathbb{R}}^{D}$ represents a single object. For scene parsing, we adopt the recursive object-centric video decomposition framework from Kipf et al. (2022).
At time step $t$, the corresponding input frame $\textbf{X}_{t}$ is encoded with a feature extractor module into a set of $D_{h}$-dimensional feature maps $\textbf{h}_{t}\in{\mathbb{R}}^{L\times D_{h}}$ representing $L$ spatial locations. The feature extractor is a convolutional neural network in our $\text{TextOCVP}_{\text{SAVi}}$ variant and a DINO-pretrained vision transformer (Caron et al., 2021) in $\text{TextOCVP}_{\text{DINO}}$. These feature maps are processed with Slot Attention (Locatello et al., 2020), which updates the previous slots $\textbf{S}_{t-1}$ based on visual features from the current frame following an iterative attention mechanism. Namely, Slot Attention performs cross-attention, with the attention weights normalized over the slot dimension, thus encouraging the slots to compete to represent parts of the input. It then updates the slots using a Gated Recurrent Unit (Cho et al., 2014) (GRU). Formally, Slot Attention updates the previous slots $\textbf{S}_{t-1}$ by:
@@ 3.2 Text-Conditioned Object-Centric Predictor
Our proposed text-conditioned predictor module, illustrated in Fig. 3, autoregressively forecasts future object states conditioned on the object slots from the reference frame $\textbf{S}_{1}$ and a text description $\mathcal{C}$. This design enables TextOCVP to predict temporally consistent object dynamics that are grounded not only in visual context, but also in high-level semantic intent specified via natural language.
To condition the prediction process, the text description $\mathcal{C}$ is encoded into text token embeddings $\mathbf{C}$ using an encoder-only transformer. These tokens provide a global semantic prior guiding predictions at every time step. We experiment with different encoder variants, including vanilla transformer encoder (Vaswani et al., 2017) and pretrained T5 (Raffel et al., 2020).
@@
[…]
Figure 3: Overview of TextOCVP’s text-conditioned object-centric video predictor transformer.
Each layer of our predictor module mirrors the transformer decoder architecture (Vaswani et al., 2017). First, a self-attention layer enables every slot to attend to all other object representations in the sequence, modeling the spatio-temporal relations between objects. Subsequently, a text-to-slot cross-attention layer enhances the slot representations by incorporating relevant features from the text embeddings, guiding the prediction process to align with the motion and dynamics described in the textual caption. Finally, an MLP is applied for each token. This process is repeated in every predictor layer, resulting in the predicted object slots of the subsequent time step $\hat{\textbf{S}}_{t+1}$. Furthermore, we apply a residual connection from $\textbf{S}_{t}$ to $\hat{\textbf{S}}_{t+1}$, which improves the prediction temporal consistency. This process is repeated autoregressively to obtain slot predictions for $T$ subsequent time steps.
@@ 3.3 Video Rendering
The video rendering module decodes the predicted slots $\hat{\textbf{S}}_{t}$ to render the corresponding video frame $\hat{\textbf{X}}_{t}$. We leverage two variants of the video rendering module, for our $\text{TextOCVP}_{\text{SAVi}}$ and $\text{TextOCVP}_{\text{DINO}}$ variants.
@@
This variant independently decodes each slot in $\hat{\textbf{S}}_{t}$ with a CNN-based Spatial Broadcast Decoder (Watters et al., 2019), rendering an object image $\textbf{o}_{t}^{n}$ and mask $\textbf{m}_{t}^{n}$ for each slot $\textbf{s}_{t}^{n}$. The object masks are normalized across the slot dimension, and the representations are combined via a weighted sum to render video frames:
@@
This decoder variant decodes the object slots in two distinct stages. First, following DINOSAUR (Seitzer et al., 2023), an MLP-based Spatial Broadcast Decoder (Watters et al., 2019) is used to generate object features along with their corresponding masks. Similar to the $\text{TextOCVP}_{\text{SAVi}}$ decoder, the object masks are normalized and combined with the object features in order to reconstruct the encoded features $\hat{\textbf{h}}_{t}\in{\mathbb{R}}^{L\times D_{h}}$. In the second stage, the reconstructed features $\hat{\textbf{h}}_{t}$ are arranged into a grid format and processed with a CNN decoder to generate the corresponding video frame $\hat{\textbf{X}}_{t}$.